OpenStudy (anonymous):

easy medal question John's race car memorabilia rose in value 3% last year. It is now worth $127.04. What was it worth a year ago?

OpenStudy (anonymous):

this is not diffcult just for fun

OpenStudy (anonymous):

Let the value last year be x. x + 3/100 * x = 127.04 Solve for x.

Parth (parthkohli):

\[\LARGE x + {3x \over 100} = 127.04\]

Parth (parthkohli):

\[\LARGE {103x \over 100} = 127.04\] \[\LARGE 103x = 12704\] Can you solve it now?

OpenStudy (anonymous):

Taking LCM we have, 103x/100 = 127.04 Transposing 103/100 to RHS, x = 127.04 * 100/103 = 123.3

OpenStudy (anonymous):

no dont explain it to me it just a easy question to see is the fastest online i alreadt now the answer

OpenStudy (anonymous):

pratu043 wins yahhhhh

OpenStudy (anonymous):

yay....

OpenStudy (anonymous):

faned and medaled

OpenStudy (anonymous):

thanks for playin parth kohli

Parth (parthkohli):

This site is about learning. Not about tests/quizzes. Please don't do this again.

OpenStudy (anonymous):

The OS mods mentioned that they might be adding a quiz feature soon.

OpenStudy (anonymous):

sorry gosh man just trying to help lighten the modd i didnt mean to make you mad for one question

OpenStudy (anonymous):

Many people are supportive of this feature.

Parth (parthkohli):

They shouldn't actually. The one who deserves the medals wouldn't get them.

OpenStudy (anonymous):

You're right, but you got a medal didn't you?

OpenStudy (anonymous):

It all good guys let just go on with our day :)

Parth (parthkohli):

That's because I explained, you too, but like in general I'm saying that in the quiz section, we'd have to do it ourself.